# Exploring Hubble Tension Alleviation through Neutrino-Coupled Perturbed $f(R)$ Gravity

### v1

@@ VIII Hubble Tension analysis
In contrast, the perturbed $f(R)$ gravity model coupled with neutrinos demonstrated a more significant potential to alleviate the Hubble tension. By incorporating the neutrino coupling, the inferred $H_{0}$ values shifted closer to those from local measurements, reducing the tension more effectively than the non-coupled case. For certain combinations of datasets, the tension with Planck 2018 decreased to $0.74\sigma$, while the tension with R22 dropped to $0.95\sigma$, highlighting the promise of this extended model. The improved performance of the neutrino-coupled model underscores its capability to address the observed discrepancies between early- and late-time cosmological measurements.
These results suggest that the perturbed $f(R)$ gravity model, particularly when coupled with neutrinos, represents a promising framework for mitigating the Hubble tension. While it does not completely resolve the discrepancy, the improved agreement with late-time measurements indicates the potential of this model to account for deviations in cosmological observations. Future work should focus on further refining this framework, including exploring additional coupling mechanisms, quantum corrections, and interactions with dark energy, to better address the unresolved tensions.
In summary, the perturbed $f(R)$ gravity model coupled with neutrinos offers a more effective approach to alleviating the Hubble tension compared to its non-coupled counterpart. This advancement paves the way for further theoretical and observational efforts to unravel the complexities of the Universe’s expansion history.

### v2

@@ VIII Hubble Tension analysis
$S_{8}$ Tension with Planck 2018, KiDS, and DES for Perturbed $f(R)$ Gravity
The parameter $S_{8}$, which combines the matter density $\Omega_{m}$ and the amplitude of matter fluctuations $\sigma_{8}$, plays a crucial role in understanding the large-scale structure of the Universe. In our analysis, we find the $S_{8}$ value for the CMB + All datasets to be $S_{8}=0.798\pm 0.030.$
This result is slightly lower than the value inferred from the Planck CMB measurement, which gives $S_{8}=0.832\pm 0.013$ Aghanim et al. (2018), indicating a moderate tension of approximately $1.43\sigma$.
Furthermore, when compared to measurements from KiDS-1000x{2dFLenS+BOSS}, which reports $S_{8}=0.766^{+0.020}_{-0.014}$ 50 , and DES-Y3, which provides $S_{8}=0.776\pm 0.017$ 51 , the tension with our CMB + All result is about $1.15\sigma$ and $1.32\sigma$, respectively.
These results underscore the ongoing challenge of reconciling cosmic microwave background data with large-scale structure observations within the framework of perturbed $f(R)$ gravity, highlighting potential discrepancies between theoretical predictions and observational data. All the above results are summarized in Table VIII.
Dataset Value Planck Tension KiDS-1000 Tension DES-Y3 Tension
CMB + Lensing $S_{8}=0.808\pm 0.066$ $0.53\sigma$ $0.49\sigma$ $0.68\sigma$
CMB + Lensing + BAO $S_{8}=0.8\pm 0.046$ $0.87\sigma$ $0.96\sigma$ $1.15\sigma$
CMB + Lensing + CC $S_{8}=0.801\pm 0.030$ $1.47\sigma$ $1.57\sigma$ $1.89\sigma$
CMB + Lensing + Pantheon $S_{8}=0.802\pm 0.048$ $0.79\sigma$ $0.81\sigma$ $1.02\sigma$
CMB + All $S_{8}=0.798\pm 0.030$ $1.43\sigma$ $1.15\sigma$ $1.32\sigma$
Table 8: Summary of $S_{8}$ tensions with Planck 2018, KiDS, and DES for the perturbed $f(R)$ gravity model.
Figure 6: The comparison of $S_{8}$ for different combination of data sets(CMB+Other) for perturbed $f(R)$ gravity model.
$S_{8}$ Tension with Planck 2018, KiDS, and DES for Perturbed $f(R)$ Gravity Coupled with Neutrinos
In our analysis within the framework of perturbed $f(R)$ gravity coupled with neutrinos, we have derived $S_{8}$ values from various combinations of observational datasets. A detailed summary of these results is presented in Table IX.
We find that the $S_{8}$ value for the CMB + All datasets is $S_{8}=0.785\pm 0.046$. When compared to the Planck 2018 result of $S_{8}=0.832\pm 0.013$ Aghanim et al. (2018), this corresponds to a tension of approximately $1.13\sigma$. This moderate discrepancy suggests a possible deviation from the standard cosmological model.
Comparisons with large-scale structure surveys reveal further tensions. The KiDS-1000x{2dFLenS+BOSS} result of $S_{8}=0.766^{+0.02}_{-0.014}$ 50 shows a tension of approximately $0.70\sigma$ with our findings, while the DES-Y3 value of $S_{8}=0.776\pm 0.017$ 51 indicates a tension of around $0.23\sigma$.
These results highlight the challenges in reconciling cosmic microwave background data with large-scale structure observations within the perturbed $f(R)$ gravity framework. The inclusion of neutrino coupling appears to reduce the level of tension with Planck and large-scale structure datasets compared to the standard cosmological model, suggesting that modified gravity and neutrino interactions may play a role in resolving the $S_{8}$ tension.
Dataset Value Tension with Planck Tension with KiDS-1000 Tension with DES-Y3
CMB + Lensing $S_{8}=0.784\pm 0.065$ $0.74\sigma$ $0.86\sigma$ $0.46\sigma$
CMB + Lensing + BAO $S_{8}=0.791\pm 0.045$ $1.13\sigma$ $0.75\sigma$ $0.34\sigma$
CMB + Lensing + CC $S_{8}=0.793\pm 0.046$ $1.06\sigma$ $0.79\sigma$ $0.39\sigma$
CMB + Lensing + Pantheon $S_{8}=0.792\pm 0.058$ $0.86\sigma$ $0.72\sigma$ $0.36\sigma$
CMB + All $S_{8}=0.785\pm 0.046$ $1.13\sigma$ $0.70\sigma$ $0.23\sigma$
Table 9: Summary of $S_{8}$ tensions with Planck 2018, KiDS-1000, and DES-Y3 for the perturbed $f(R)$ gravity model coupled with neutrinos.
The results show that the level of tension with the Planck 2018 results remains significant but is somewhat reduced when compared to the standard cosmological model. Interestingly, the tensions with KiDS-1000 and DES-Y3 are relatively low, suggesting that the perturbed $f(R)$ gravity model with neutrino coupling may provide a better fit to large-scale structure data.
Figure 7: The comparison of $S_{8}$ for different combination of data sets(CMB+Other) for perturbed $f(R)$ gravity model coupled with neutrinos.
Figures 6 and 7 present a comparison of the clustering amplitude parameter $S_{8}$ for different combinations of data sets, including the cosmic microwave background (CMB) and other observational probes, within the context of perturbed $f(R)$ gravity models. Figure 6 shows the results for the perturbed $f(R)$ gravity model coupled without neutrinos, highlighting the impact of neutrino coupling on the clustering amplitude. In contrast, Figure 7 illustrates the comparison for the perturbed $f(R)$ gravity model with neutrino coupling, providing insight into the effects of modified gravity alone on the $S_{8}$ parameter. These figures underscore the role of neutrino interactions and modified gravity in resolving the observed tension in the clustering amplitude.
@@
In contrast, the perturbed $f(R)$ gravity model coupled with neutrinos demonstrated a more significant potential to alleviate the Hubble tension. By incorporating the neutrino coupling, the inferred $H_{0}$ values shifted closer to those from local measurements, reducing the tension more effectively than the non-coupled case. For certain combinations of datasets, the tension with Planck 2018 decreased to $0.74\sigma$, while the tension with R22 dropped to $0.95\sigma$, highlighting the promise of this extended model. The improved performance of the neutrino-coupled model underscores its capability to address the observed discrepancies between early- and late-time cosmological measurements. Our analysis of the parameter $S_{8}$, which combines the matter density $\Omega_{m}$ and the amplitude of matter fluctuations $\sigma_{8}$, reveals moderate tension between the values inferred from cosmic microwave background (CMB) data and large-scale structure surveys within the framework of perturbed $f(R)$ gravity models, both with and without neutrino coupling. For the non-interacting case, the $S_{8}$ value for the CMB + All dataset is $S_{8}=0.798\pm 0.030$, showing a $1.43\sigma$ tension with Planck 2018 and tensions of approximately $1.15\sigma$ and $1.32\sigma$ with KiDS-1000 and DES-Y3, respectively. For the interacting case with neutrino coupling, the CMB + All result is $S_{8}=0.785\pm 0.046$, reducing the tension with Planck to $1.13\sigma$ and lowering the tensions with KiDS-1000 and DES-Y3 to $0.70\sigma$ and $0.23\sigma$, respectively. These findings suggest that modified gravity and neutrino interactions may help mitigate the $S_{8}$ tension and improve consistency with large-scale structure data. These results suggest that the perturbed $f(R)$ gravity model, particularly when coupled with neutrinos, represents a promising framework for mitigating the Hubble and $S_{8}$ tension. While it does not completely resolve the discrepancy, the improved agreement with late-time measurements indicates the potential of this model to account for deviations in cosmological observations. Future work should focus on further refining this framework, including exploring additional coupling mechanisms, quantum corrections, and interactions with dark energy, to better address the unresolved tensions.
In summary, the perturbed $f(R)$ gravity model coupled with neutrinos offers a more effective approach to alleviating the Hubble and $S_{8}$ tension compared to its non-coupled counterpart. This advancement paves the way for further theoretical and observational efforts to unravel the complexities of the Universe’s expansion history.